how would this story problem work?

Char inherited $30,000 from her great uncle. She invested it in 2 different funds. One earns 5% adn the other earns 6%. If she earned $1600 last year, how much did she invest in each fund?

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

Char inherited $30,000 from her great uncle. She invested it in 2 different funds. One earns 5% adn the other earns 6%. If she earned $1600 last year, how much did she invest in each fund?
------------
Let amt. invested at 5% by "x"; interest on this is 0.05x dollars
--------------
Amt. invested at 6% is "30,000-x" ; int is 0.06(30000-x)=1800-0.06x dollars
---------------------
EQUATION:
interest + interest = 1600 dollars
0.05x + 1800-0.06x = 1600
-0.01x = -200
x = $20,000 (amt. invested at 5%)
------------
30,000-x = $10,000 (amt. invested at 6%)
===============
Cheers,
Stan H.

Answer by checkley71(8403)

.06X+.05(30,000-X)=1,600
.06X+1500-.05X=1,600
.06X-.05X=1,600-1,500
.01X=100
X=100/.01
X=10,000 INVESTED @ 6%.
30,000-10,000=20,000 INVESTED @ 6%.
PROOF
.06*10,000+.05*20,000=1,600
600+1,000=1,600
1,600=1,600